Find the derivative of the given function.


1.) y = (x ^ 3 - 3)(x ^ 2 + 4x + 1)


2.) y = (7x + 3)(x ^ 4 - x ^ 3 - 9x)


3.) y = x ^ 2 * (2x ^ 2 + x + 1)


4.) y = (x ^ 2 - 10x + 2)(x ^ 3 - 2x ^ 2 + 1)


5.) y=(x^ 2 -2x-1)(x+)


6.) y= sqrt x( x^ 3 -2x^ 2 +7x-1)


7.) t = x ^ (1/3) * (x ^ 2 - 5x + 2)


8.) y = (5x ^ 2 - 13)(x ^ 3 + 6x + 1)


9.) y=x^ -2 (x^ 2 +7


10.) x ^ - 5 * (x ^ 2 + 10x - 5)


1
Expert's answer
2022-03-25T05:30:16-0400

"1. y=(x^3-3)(x^2+4x+1)\\
y'=3x^2(x^2+4x+1)+(x^3-3)(2x+4)=\\=5x^4+16x^3+3x^2-6x-12\\
2. y=(7x+3)(x^4-x^3-9x)\\
y'=7(x^4-x^3-9x)+(7x+3)(4x^3-3x^2-9)=\\
=35x^4-16x^3-9x^2-126x-27\\
3. y=x^2(2x^2+x+1)\\
y'=2x(2x^2+x+1)+x^2(4x+1)=\\
=8x^3+3x^2+2x\\
4. y=(x^2-10x+2)(x^3-2x^2+1)\\
y'=(2x-10)(x^3-2x^2+1)+\\+(x^2-10x+2)(3x^2-4x)=\\
=5x^4-48x^3+66x^2-6x-10\\
5. y=(x^2-2x-1)(x+1)\\
y'=(2x-2)(x+1)+(x^2-2x-1)=\\
=3x^2-2x-3\\
6. y=\sqrt{x^3-2x^2+7x-1}\\
y'=\frac{1}{2\sqrt{x^3-2x^2+7x-1}}(3x^2-4x+7)\\
7. t=x^{\frac{1}{3}}(x^2-5x+2)\\
t'=\frac{1}{3}x^{-\frac{2}{3}}(x^2-5x+2)+x^{\frac{1}{3}}(2x-5)\\
8. y=(5x^2-13)(x^3+6x+1)\\
y'=10x(x^3+6x+1)+(5x^2-13)(3x^2+6)=\\
=25x^4+51x^2+10x-78\\
9. y=x^{-2}(x^2+7)\\
y'=-2x^{-3}(x^2+7)+x^{-2}\cdot2x\\10. 
y=x^{-5}(x^2+10x-5)\\
y'=-5x^{-6}(x^2+10x-5)+x^{-5}(2x+10)"